OpenStudy (anonymous):

Find two unit vectors orthogonal to $$\vec{a}=<−5,3,2>$$ and $$\vec{b}=<3,−4,−4>$$

OpenStudy (unklerhaukus):

say one of the vectors you are looking for is \[\vec c = \langle c_1,c_2,c_3\rangle\] For orthogonality \[\vec a\cdot\vec c=0\] i.e. \[\langle -5,3,2\rangle\cdot\langle c_1,c_2,c_3\rangle=0\]

OpenStudy (unklerhaukus):

compute this dot product, and do the same for the vector \(\vec b\)

OpenStudy (unklerhaukus):

you should get two equations in-terms of \(c_1\), \(c_2\), and \(c_3\)

OpenStudy (anonymous):

doesn't c need to be orthogonal to both a and b?

OpenStudy (unklerhaukus):

yes, the two equations you get should both be true

OpenStudy (anonymous):

im attempting this now.. not sure i know what im doing though

OpenStudy (anonymous):

i mean I'm attempting what you said, im very slow though

OpenStudy (unklerhaukus):

*\[\langle -5,3,2\rangle\cdot\langle c_1,c_2,c_3\rangle=0\\-5c_1+3c_2+2c_3=0\]

OpenStudy (anonymous):

ok so the other one is $$3c_{1} -4c_{2} - 4c_{2}$$

OpenStudy (anonymous):

= 0

OpenStudy (unklerhaukus):

yep

OpenStudy (unklerhaukus):

So the components of \(\vec c\) , must satisfy both \[-5c_1+3c_2+2c_3=0\]and\[3c_1-4c_2-4c_3\]

OpenStudy (anonymous):

set them equal to each other now since they both equal 0?

OpenStudy (unklerhaukus):

eliminate one of the components and solve for another component in-terms of the other one

OpenStudy (anonymous):

so i can multiply the top one by -2 right?

OpenStudy (anonymous):

to get rid of c3?

OpenStudy (unklerhaukus):

yes! that is a good idea

OpenStudy (anonymous):

awesome, I'm going to try this, ty so much :) I think i can get it now, i hope :) lol

OpenStudy (anonymous):

@UnkleRhaukus, still not getting it, ty for the help though, is there another way to do this using the cross product?

OpenStudy (anonymous):

I'm getting c1 = 2a/7 c2 = a and c3 = -11a/14

OpenStudy (anonymous):

Ah, I got it was not reading question correctly, I needed to find unit vectors, ty again :)